You have 15 pennies in your pocket. Of those pennies, 3 are Canadian. Suppose you pick a penny out of your pocket at random. Find P(not Canadian)
this your answer hun
P ( canadian)= 3/15 = 1/5
-> P(not canadian) = 1 - 1/5 = 4/5
Answer:
P(Not Canadian) = 0.8
Step-by-step explanation:
Be the events
C = the penny is Canadian
NC = the penny is NOT Canadian
The probability of an event is equal to the successful events among the total events.
\(P(C)=\frac{3 (Canadian Cents)}{15(Cents in total)}=\frac{3}{15}=0.2\)
The sum of the total probability is 1. Then,
P (NC) = 1 - PC = 1 - 0.2 = 0.8
The probability that a random penny is non-Canadian is 0.8
Hope this helps!